if f(x)= 2x-3, then compute f(f(f(f(f(f(3))))))
Answer by Edwin McCravy(6929)


if f(x)= 2x-3, then compute f(f(f(f(f(f(3))))))

f(3) = 2(3) - 3 = 3

so that means 

f(f(3)) = f(3) = 3 

and

f(f(f(3))) = f(3) = 3

and

f(f(f(f(3)))) = f(3) = 3

...

and you can reason from there that you'll always get 3.

So the answer is 3.
 
Edwin